# Multi-scale analysis of flow over heterogeneous urban environments

### v1

@@ 4.3 Drag distribution
As discussed in §2.4, the drag distribution $\left\langle{\smash{f_{D}}}\right\rangle$ or equivalently the cumulative kinematic shear stress $\left\langle{\smash{\tau_{D}}}\right\rangle$ is a key quantity that requires parameterisation for non-building resolving NWP simulation. Figure 9 shows the plane-averaged value of the cumulative drag force $\left\langle{\smash{\tau_{D;L}}}\right\rangle_{xy}$ at different averaging lengths $L$ together with its standard deviation $\sigma_{R}$ as a function of the scaled height $1-\zeta$. Here, we have chosen $1-\zeta$ over $\zeta$ so the canopy top is presented at the top of the figures. Note that $\left\langle{\smash{\tau_{D;L}}}\right\rangle_{xy}$ is independent of $L$, as shown in the black line. The variances are typically small near the canopy top and increase with decreasing height. It is also clearly visible that the spread in $\tau_{D;L}$ increases as the averaging length $L$ decreases. This suggests that the plane-averaged drag is not representative of the local value at a specific height.
Below we test the performance of the Sützl et al. (2020) parameterisation (36) at different averaging lengths $L$. In order to do this, we need to make the parameterisation local, which can be done as follows:
@@
$\zeta_{L}(\boldsymbol{x}_{\perp},z)=\frac{1}{\lambda_{f;L}}\int_{z}^{h}\left\langle{\smash{\rho_{L}}}\right\rangle_{L}(\boldsymbol{x}_{\perp},z^{\prime})\,\textnormal{d}z^{\prime}\,,$ (41)
@@
$\lambda_{f;L}(\boldsymbol{x}_{\perp})=\int_{0}^{h}\left\langle{\smash{\rho_{L}}}\right\rangle_{L}(\boldsymbol{x}_{\perp},z^{\prime})\,\textnormal{d}z^{\prime}\,.$ (42)
@@
Figure 10: Normalised local cumulative drag stress $\tau_{D;L}/\tau_{0;L}$ against the local scaled frontal area $\zeta_{L}$ at different averaging lengths. (a) $L=96$ m, (b) $L=192$ m, (c) $L=384$ m, (d) $L=786$ m. The dashed line represents the parametrisation (40). The coefficient of determination $R^{2}$ is labelled to show the performance of parametrisation.
@@
Figure 10 shows the distribution density of the local cumulative drag stress $\tau_{D;L}$ against the local scaled frontal area $\zeta_{L}$ at different resolutions overlaid with the parametrisation (40). Figures 10(c,d) show that at averaging lengths $L=384$ and $786$ m, there is very little scatter in the data as the fields are relatively homogeneous and the parametrisation works well with a very high coefficient of determination. This clearly shows that accurate parameterisation requires local quantities rather than plane-averaged quantities (Figure 9).
@@
As the resolution increases, the scatter increases due to heterogeneity effects. Indeed, (20) shows that, whenever there are heterogeneity effects, the horizontal terms in the momentum equation become important, which implies that the vertical terms are no longer in balance. The latter is of course what (40) is built on. However, a high coefficient of determination $R^{2}$ illustrates that the parameterisation still fits the data reasonably well at $L=96$ m (figure 10a). However, at averaging lengths below 96 m, the agreement between the data and the parameterisation quickly deteriorates, we find that the coefficient of determination dramatically reduces to $0.23$ at $L=48$ m which means a poor match. Indeed, below $L=96$ m, there is a significant increase in the resolved variance (figure 8), and thus the resolved heterogeneity. Note that, the data distribution within this figure is much narrower than that shown in figure 6 which is associated with the same quantity $\tau_{D;L}$, the reason is that buildings with different heights would accumulate various of drag stress from their bottom to top, i.e., at $\zeta=0$, therefore, figure 6 shows a wide deviation. However, in figures 10, normalising the cumulative stress by its local value greatly reduces the spread of the data.
@@
Figure 11: Comparison between the cumulative drag $\tau_{0;L}$ and the shear stress parameterisation (43) at various averaging lenghts. (a) $L=96$ m, (b) $L=192$ m, (c) $L=384$ m, (d) $L=786$ m. The dashed line indicates perfect agreement. The coefficient of determination $R^{2}$ is labelled to show the performance of parametrisation.
@@
A potential drawback of the parameterisation (40) is that, even though it agrees with the data well at averaging lengths larger than 96 m, it does not provide the information for the superficially averaged local shear stress $\tau_{0;L}$. Here we model $\tau_{0;L}$ using a single drag coefficient $\hat{C}_{d}$ as (Coceal and Belcher 2004; Belcher 2005; Sützl et al. 2021)
@@
where $U=\left\langle{\smash{\overline{u}}}\right\rangle_{L}(\boldsymbol{x}_{\perp},z=30)$ is the reference (superficial) wind velocity at $30$ m which coincides with the mean building height. Note that $\hat{C}_{d}$ is distinct from $C_{d}$ in figure 6, which is height-dependent. Using the plane-averaged data, the appropriate value for $\hat{C}_{d}$ in the parameterisation above is $\hat{C}_{d}=2.49$.
In Figure 11 we show the distribution of $\tau_{0;L}$ against the parametrisation $\frac{1}{2}\hat{C}_{d}\lambda_{f;L}U^{2}$. For $L=768$ m (figure 11d), the data has a very small spread and has the right value, as expected since $\hat{C}_{d}$ was inferred from the plane-averaged data. As $L$ becomes smaller, the spread of the data becomes wider, although still centred around the diagonal. For $L=96$ m (figure 11a), the parametrisation does not predict the local surface stress well. However, the coefficients of determination $R^{2}$ of all four figures are very poor, which is because $R^{2}$ becomes less informative when the data show little variation or are very close to their mean value, even for good modellings. With this regard, we also evaluate the normalised root-mean-square error (NRMSE) to identify the performance of the parametrisation — the lower the NRMSE is, the better the model predicts. This error is generally consistent with the data distribution in the figure.
@@ 4.4 Canopy turbulence closures
As discussed in §2.4, the other key process that requires parameterisation for non-building resolving NWP simulation is the unresolved stress $\left\langle{\smash{\tau_{i}}}\right\rangle$. Consistent with (17,18), we consider the vertical flux $\left\langle{\smash{\tau_{z}}}\right\rangle=\left\langle{\smash{\overline{w}}}\right\rangle\left\langle{\smash{\overline{u}}}\right\rangle-\left\langle{\smash{\overline{w}\,\overline{u}}}\right\rangle-\left\langle{\smash{\overline{w^{\prime}u^{\prime}}}}\right\rangle$, which comprises a turbulent momentum flux $\left\langle{\smash{\overline{w^{\prime}u^{\prime}}}}\right\rangle$ and a residual term $\left\langle{\smash{\overline{w}}}\right\rangle\left\langle{\smash{\overline{u}}}\right\rangle-\left\langle{\smash{\overline{w}\,\overline{u}}}\right\rangle$. As demonstrated in (12), this residual term can be further decomposed associated, but for simplicity, we regard the residual as a whole and name it a dispersive-like term.
@@
The eddy-diffusivity coefficient $K_{m}$ can be modelled in many different ways (Hanjalić and Launder 2022). We investigate two models here, the $k-l$ model (Deardorff 1972; Bougeault and Lacarrère 1989) and the $k-\omega$ model (Wilcox 1998):
@@
Figure 12: Suitability of various turbulence closures to parameterise the turbulence shear stress $-\left\langle{\smash{\overline{w^{\prime}u^{\prime}}}}\right\rangle_{xy}$ (top row) and the total unresolved stress $\left\langle{\smash{\tau_{z}}}\right\rangle_{xy}$ (bottom row). (a,e) Constant eddy viscosity, (b,f) $k-l$ model, (c,g) $k-\omega$ model, and (d,h) parameterisation (40). Agreement of the data with the dashed lines implies agreement with the model. The data are shown within the canopy region. The coefficient of determination $R^{2}$ is labelled to show the performance of parametrisation.
@@
First, we will study the closures for plane-averaged statistics. Figures 12(a-c) show the turbulence fluxes $\left\langle{\smash{\overline{w^{\prime}u^{\prime}}}}\right\rangle_{xy}$ in the canopy layer with the x-axes chosen such that a straight line starting at the origin indicates agreement with the closure. Figure 12(a) shows that the turbulence flux in the canopy region is practically independent of $S_{31}$, suggesting that the eddy diffusivity $K_{m}$ cannot be assumed constant in the canopy layer. The best-fit value (associated with the dashed line is $K_{m}=8.91$ m2/s, which is greater than the value of $1$ m2/s reported by Lu et al. (2024). This is likely due to the difference in geometry. However, the negative coefficient of determination $R^{2}$ indicates that the line fitting from the origin shows a poor match. Figure 12(b) shows that the mixing length model performs better, but although the data is pretty much on a straight line it does not pass through the origin. The best-fit results in $l_{k}=13.00$ m, here we have assumed the model constant $C_{k}=0.4$ (Bougeault and Lacarrère 1989). The ratio $C_{k}l_{k}/h_{\max}=0.07$ is comparable with (Nazarian et al. 2020) at the same urban density $\lambda_{p}$. Only for the $k-\omega$ model (Figure 12c) does the data compare well with a line through the origin (with $R^{2}$ very close to $1$). The best-fit returns a value $\omega=0.40$ s-1.
@@
Figures 12(e-g) present the same figures, but now using the kinematic unresolved shear stress $\left\langle{\smash{\tau_{z}}}\right\rangle_{xy}$ which includes the dispersive-like term. The small difference between Figures 12(a-c) and (e-g) is due to the fact that the turbulence fluxes dominate everywhere in the canopy layer (Figure 4b). All the figures show that considering the dispersive-like terms slightly reduces the models’ performance. However, again, the $k-\omega$ model can be seen to perform the best with the best-fit values $\omega=0.33$ s-1.
@@
Another way to parameterise $\left\langle{\smash{\tau_{z}}}\right\rangle$ without having to rely on auxiliary variables like $k$ is to use the drag parametrisation (40). Invoking the constant stress layer assumption often used in boundary layer theory (Schlichting and Gersten 2017) implies that $\tau_{f}\approx\tau_{0}$, which after substitution into (25) and making use of (40) immediately results in
@@
The distribution of $\left\langle{\smash{\overline{w^{\prime}u^{\prime}}}}\right\rangle_{xy}$ and $\left\langle{\smash{\tau_{z}}}\right\rangle_{xy}$ against $\zeta$ are shown in figures 12(d,h), respectively. Since the parameterisation is derived for $\left\langle{\smash{\tau_{z}}}\right\rangle_{xy}$, it always overestimates the $\left\langle{\smash{\overline{w^{\prime}u^{\prime}}}}\right\rangle_{xy}$ by definition — the total stress is in better agreement with the parameterisation than the turbulence flux only (figure 12h). Near the top, for $0.91-\zeta>$ 0.9, the parametrisation slightly overestimates the total stress. Comparing the bottom row, the parametrisations in figure 12(g,h) of total unresolved stress have the largest $R^{2}$ to fit the data.
Figure 13: Skill of the $k-\omega$ model (top row) and (40) (bottom row) to parameterise the total unresolved vertical stress $\left\langle{\smash{\tau_{z}}}\right\rangle_{L}$ at various averaging lengths. (a,e) $L=96$ m, (b,f) $L=192$ m, (c,g) $L=384$ m, (d,h) $L=768$ m. Overlaid with the parametrisation from figure 12(g,h) for the top and bottom row, respectively. The coefficient of determination $R^{2}$ is labelled to show the performance of parametrisation.
Figure 13 shows the two best-performing parameterisations of total unresolved stress at different averaging lengths $L$. The top panels show the $k-\omega$ model predictions for the total unresolved shear stress $\left\langle{\smash{\tau_{z}}}\right\rangle_{L}$ and the bottom row shows the predictions derived from the drag parameterisation (46). Note that the stresses in the bottom row are normalised by $\left\langle{\smash{\tau_{z}}}\right\rangle_{L}$ at maximum building height $z=h_{\max}$ in order to converge to the unit. However, under the constant stress layer assumption (Schlichting and Gersten 2017), $\left\langle{\smash{\tau_{z}}}\right\rangle_{L}(z=h_{\max})\approx\tau_{0}$ which is consistent with the parameterisation (46). In order to create a scale-dependent parameterisation for (46), simply replace $\left\langle{\smash{\tau_{z}}}\right\rangle_{xy}$ with $\left\langle{\smash{\tau_{z}}}\right\rangle_{L}$, $\tau_{0}$ with $\tau_{0;L}$, and $\zeta$ with $\zeta_{L}$. As expected, as the averaging length reduces, the data spread around the diagnostic line increases with the reduced coefficient of determination $R^{2}$. However, $R^{2}$ is larger in the bottom row than the top row at the same resolution, suggesting that the parametrisation (46) has higher fidelity than the $k-\omega$ closure. We also note that, compared with the drag parametrisation (figure 10), the coefficients $R^{2}$ of stress parametrisation (46) are lower at the same resolution, and reduce more rapidly with the reducing resolution. This might be because of the assumption of constant stress.

### v2

@@ 4.3 Drag distribution
As discussed in §2.4, the drag distribution $\left\langle{\smash{f_{D}}}\right\rangle$ or equivalently the cumulative kinematic shear stress $\left\langle{\smash{\tau_{D}}}\right\rangle$ is a key quantity that requires parameterisation for non-building resolving NWP simulations. Figure 9 shows the plane-averaged value of the cumulative drag force $\left\langle{\smash{\tau_{D;L}}}\right\rangle_{xy}$ at different averaging lengths $L$ together with its standard deviation $\sigma_{R}$ as a function of the scaled height $1-\zeta$. Here, we have chosen $1-\zeta$ over $\zeta$ so the canopy top is presented at the top of the figures. Note that $\left\langle{\smash{\tau_{D;L}}}\right\rangle_{xy}$ is independent of $L$, as shown in the black line. The variances are typically small near the canopy top and increase with decreasing height. It is also clearly visible that the spread in $\tau_{D;L}$ increases as the averaging length $L$ decreases. This suggests that the plane-averaged drag is not representative of the local value at a specific height.
In the following, we perform an a priori analysis at different averaging lengths $L$ and test the performance of the Sützl et al. (2020) parameterisation Eq. (36). In order to do this, we need to make the parameterisation local, which can be done as follows:
@@
$\zeta_{L}(\boldsymbol{x}_{h},z)\equiv\frac{1}{\lambda_{f;L}}\int_{z}^{h}\left\langle{\smash{\rho_{L}}}\right\rangle_{L}(\boldsymbol{x}_{h},z^{\prime})\,\textnormal{d}z^{\prime}\,$, (41)
@@
$\lambda_{f;L}(\boldsymbol{x}_{h})\equiv\int_{0}^{h}\left\langle{\smash{\rho_{L}}}\right\rangle_{L}(\boldsymbol{x}_{h},z^{\prime})\,\textnormal{d}z^{\prime}\,.$ (42)
@@
Figure 10: Normalised local cumulative drag stress $\tau_{D;L}/\tau_{0;L}$ against the local scaled frontal area $\zeta_{L}$ at different averaging lengths. The colour represents the statistical probability density function value of the data appearing at each location. (a) $L=96$ m, (b) $L=192$ m, (c) $L=384$ m, (d) $L=786$ m. The dashed line represents the parametrisation Eq. (40). The coefficient of determination $R^{2}$ is labelled to show the a priori performance of parametrisation.
@@
Figure 10 shows the probability density function (PDF) value of the distribution of the local cumulative drag stress $\tau_{D;L}$ against the local scaled frontal area $\zeta_{L}$, at different resolutions and overlaid with the parametrisation Eq. (40). Figures 10(c,d) show that at averaging lengths $L=384$ and $786$ m, there is very little scatter in the data as the fields are relatively homogeneous and the parametrisation works well with a very high coefficient of determination. This clearly shows that accurate parameterisation requires local quantities rather than plane-averaged quantities (Fig. 9).
@@
As the resolution increases, the scatter increases due to heterogeneity effects. Indeed, Eq. (20) shows that, whenever there are heterogeneity effects, the horizontal terms in the momentum equation become important, which implies that the vertical terms are no longer in balance. The latter is, of course, what Eq. (40) is built on. However, a high coefficient of determination $R^{2}$ illustrates that the parameterisation still fits the data reasonably well at $L=96$ m (Fig. 10a), indicating that the vertical terms still dominate. However, at averaging lengths below 96 m, the agreement between the data and the parameterisation quickly deteriorates, we find that the coefficient of determination dramatically reduces to $0.23$ at $L=48$ m, which means a poor match, indicating that further decreasing $L$, the horizontal terms can no longer be ignored. Indeed, below $L=96$ m, there is a significant increase in the resolved variance (Fig. 8), and thus the resolved heterogeneity. Note that, the data distribution within Fig. 10 is much narrower than that shown in Fig. 9 which is associated with the same quantity $\tau_{D;L}$, the reason is that buildings with different heights would accumulate various of drag stress from their bottom to top, therefore, Fig. 9 shows a wide deviation. However, in Fig. 10, normalising the cumulative stress by its local value greatly reduces the spread of the data.
@@
Figure 11: Comparison between the cumulative drag $\tau_{0;L}$ and the shear stress parameterisation Eq. (43) at various averaging lenghts. (a) $L=96$ m, (b) $L=192$ m, (c) $L=384$ m, (d) $L=786$ m. The dotted line indicates perfect agreement. The coefficient of determination $R^{2}$ is labelled to show the performance of the parametrisation.
@@
A potential drawback of Eq. (40) is that, even though it agrees with the data well at averaging lengths larger than 96 m, it does not provide the information for the superficially averaged local shear stress $\tau_{0;L}$. Here we model $\tau_{0;L}$ using a single drag coefficient $\hat{C}_{d}$ as (Coceal and Belcher 2004; Belcher 2005; Sützl et al. 2021)
@@
where $U=\left\langle{\smash{\overline{u}}}\right\rangle_{L}(\boldsymbol{x}_{h},z=30)$ is the reference (superficial) wind velocity at $30$ m which coincides with the mean building height. Note that $\hat{C}_{d}$ is distinct from $C_{d}$ in Fig. 6, which is height-dependent. Using the plane-averaged data, the appropriate value for $\hat{C}_{d}$ in the parameterisation above is $\hat{C}_{d}=2.49$.
In Fig. 11 we show the distribution of $\tau_{0;L}$ against the parametrisation $\frac{1}{2}\hat{C}_{d}\lambda_{f;L}U^{2}$. For $L=768$ m (Fig. 11d), the data has a very small spread and has the right value, as expected since $\hat{C}_{d}$ was inferred from the plane-averaged data. As $L$ becomes smaller, the spread of the data becomes wider, although still centred around the diagonal. For $L=96$ m (Fig. 11a), the parametrisation does not predict the local surface stress well. However, the coefficients of determination $R^{2}$ of all four figures are very poor, which is because $R^{2}$ becomes less informative when the data show little variation or are very close to their mean value, even for good modellings. With this regard, we also evaluate the normalised root-mean-square error (NRMSE) to identify the a priori performance of the parametrisation — the lower the NRMSE is, the better the model predicts. This error is generally consistent with the data distribution in the figure.
@@ 4.4 Canopy turbulence closures
As discussed in §2.4, the other key process that requires parameterisation for non-building resolving NWP simulation is the unresolved stress $\left\langle{\smash{\tau_{i}}}\right\rangle$. Consistent with (17,18), we consider the vertical flux $\left\langle{\smash{\tau_{z}}}\right\rangle=\left\langle{\smash{\overline{w}}}\right\rangle\left\langle{\smash{\overline{u}}}\right\rangle-\left\langle{\smash{\overline{w}\,\overline{u}}}\right\rangle-\left\langle{\smash{\overline{w^{\prime}u^{\prime}}}}\right\rangle$, which comprises a turbulent momentum flux $\left\langle{\smash{\overline{w^{\prime}u^{\prime}}}}\right\rangle$ and a residual term $\left\langle{\smash{\overline{w}}}\right\rangle\left\langle{\smash{\overline{u}}}\right\rangle-\left\langle{\smash{\overline{w}\,\overline{u}}}\right\rangle$. As demonstrated in Eq. (12), this residual term can be further decomposed associated, but for simplicity, we regard the residual as a whole and name it a dispersive-like term.
@@
The eddy-diffusivity coefficient $K_{m}$ can be modelled in many different ways (Hanjalić and Launder 2022). We investigate two models here, the klk-l model (Deardorff 1972; Bougeault and Lacarrère 1989) and the $k-\omega$ model (Wilcox 1998):
@@
Figure 12: Suitability of various turbulence closures to parameterise the turbulence shear stress $\left\langle{\smash{\overline{w^{\prime}u^{\prime}}}}\right\rangle_{xy}$ (top row) and the total unresolved stress $\left\langle{\smash{\tau_{z}}}\right\rangle_{xy}$ (bottom row). (a,e) Constant eddy viscosity, (b,f) klk-l model, (c,g) $k-\omega$ model, and (d,h) parameterisation Eq. (40). Agreement of the data with the dashed lines implies agreement with the model. The data are shown within the canopy region. The coefficient of determination $R^{2}$ is labelled to show the a priori performance of parametrisation.
@@
First, we will study the closures for plane-averaged statistics. Figures 12(a-c) show the turbulence fluxes $\left\langle{\smash{\overline{w^{\prime}u^{\prime}}}}\right\rangle_{xy}$ in the canopy layer with the x-axes chosen such that a straight line starting at the origin indicates agreement with the closure. Figure 12(a) shows that the turbulence flux in the canopy region is practically independent of $S_{31}$, suggesting that the eddy diffusivity $K_{m}$ cannot be assumed constant in the canopy layer. The best-fit value (associated with the dashed line is $K_{m}=8.91$ m2/s, which is greater than the value of $1$ m2/s reported by Lu et al. (2024). This is likely due to the difference in geometry. However, the negative coefficient of determination $R^{2}$ indicates that the line fitting from the origin shows a poor match. Figure 12(b) shows that the mixing length model performs better, but although the data is pretty much on a straight line it does not pass through the origin. The best-fit results in $l_{k}=13.00$ m, here we have assumed the model constant $C_{k}=0.4$ (Bougeault and Lacarrère 1989). The ratio $C_{k}l_{k}/h_{\max}=0.07$ is comparable with (Nazarian et al. 2020) at the same urban density $\lambda_{p}$. Only for the $k-\omega$ model (Fig. 12c) does the data compare well with a line through the origin (with $R^{2}$ very close to $1$). The best-fit returns a value $\omega=0.40$ s-1.
@@
Figures 12(e-g) present the same figures, but now using the kinematic unresolved shear stress $\left\langle{\smash{\tau_{z}}}\right\rangle_{xy}$ which includes the dispersive-like term. The small difference between Figures 12(a-c) and (e-g) is due to the fact that the turbulence fluxes dominate everywhere in the canopy layer (Fig. 4b). All the figures show that considering the dispersive-like terms slightly reduces the models’ performance. However, again, the $k-\omega$ model can be seen to perform the best with the best-fit values $\omega=0.33$ s-1.
@@
Another way to parameterise $\left\langle{\smash{\tau_{z}}}\right\rangle$ without having to rely on auxiliary variables like $k$ is to use the drag parametrisation Eq. (40). Invoking the constant stress layer assumption often used in boundary layer theory (Schlichting and Gersten 2017) implies that $\tau_{f}\approx\tau_{0}$ in the canopy layer, which after substitution into Eq. (25) and making use of Eq. (40) immediately results in
@@
The distribution of $\left\langle{\smash{\overline{w^{\prime}u^{\prime}}}}\right\rangle_{xy}$ and $\left\langle{\smash{\tau_{z}}}\right\rangle_{xy}$ against $\zeta$ are shown in figures 12(d,h), respectively. Since the parameterisation is derived for $\left\langle{\smash{\tau_{z}}}\right\rangle_{xy}$, it always overestimates the $\left\langle{\smash{\overline{w^{\prime}u^{\prime}}}}\right\rangle_{xy}$ by definition — the total stress is in better agreement with the parameterisation than the turbulence flux only (Fig. 12h). Near the top, for $1-\zeta>0.9$, the parametrisation slightly overestimates the total stress. Comparing the bottom row, the parametrisations in Fig. 12(g,h) of total unresolved stress have the largest $R^{2}$ to fit the data.
Figure 13: Skill of the $k-\omega$ model (top row) and Eq. (40) (bottom row) to parameterise the total unresolved vertical stress $\left\langle{\smash{\tau_{z}}}\right\rangle_{L}$ at various averaging lengths. (a,e) $L=96$ m, (b,f) $L=192$ m, (c,g) $L=384$ m, (d,h) $L=768$ m. Overlaid with the parametrisation from Fig. 12(g,h) for the top and bottom row, respectively. The coefficient of determination $R^{2}$ is labelled to show the a priori performance of parametrisation.
Figure 13 shows the two best-performing parameterisations of total unresolved stress at different averaging lengths $L$. The top panels show the $k-\omega$ model predictions for the total unresolved shear stress $\left\langle{\smash{\tau_{z}}}\right\rangle_{L}$ and the bottom row shows the predictions derived from the drag parameterisation Eq. (46). Note that the stresses in the bottom row are normalised by $\left\langle{\smash{\tau_{z}}}\right\rangle_{L}$ at maximum building height $z=h_{\max}$ in order to converge to the unit. However, under the constant stress layer assumption (Schlichting and Gersten 2017), $\left\langle{\smash{\tau_{z}}}\right\rangle_{L}(z=h_{\max})\approx\tau_{0}$ which is consistent with the parameterisation Eq. (46). In order to create a scale-dependent parameterisation for Eq. (46), simply replace $\left\langle{\smash{\tau_{z}}}\right\rangle_{xy}$ with $\left\langle{\smash{\tau_{z}}}\right\rangle_{L}$, $\tau_{0}$ with $\tau_{0;L}$, and $\zeta$ with $\zeta_{L}$. As expected, as the averaging length reduces, the data spread around the diagnostic line increases with the reduced coefficient of determination $R^{2}$. However, $R^{2}$ is larger in the bottom row than the top row at the same resolution, suggesting that the parametrisation Eq. (46) has higher fidelity than the $k-\omega$ closure. We also note that, compared with the drag parametrisation (Fig. 10), the coefficients $R^{2}$ of stress parametrisation Eq. (46) are lower at the same resolution, and reduce more rapidly with the reducing resolution. This might be because of the assumption of constant stress.